Hypothesis Test. Are young men more likely to experience ADD/ADHD symptoms than young women? A recent survey found 22 of 78 females had experienced ADD/ADHD symptoms while 12 of 40 males had experienced them. Test at the .1 level.

Question: Identity the correct null hypothesis.

Question 1 options:

H0:pM>pFH0:pM>pF H0:pM<pFH0:pM<pF Ha:pM=pFHa:pM=pF H0:pM=pFH0:pM=pF

None of these.

Question 2 (1 point)

Saved

Hypothesis Test. Are young men more likely to experience ADD/ADHD symptoms than young women? A recent survey found 22 of 78 females had experienced ADD/ADHD symptoms while 12 of 40 males had experienced them. Test at the .1 level.

Question: Identity the correct alternate hypothesis.

Question 2 options:

Ha:pF=pMHa:pF=pM Ha:pF<pMHa:pF<pM Ha:pF>pMHa:pF>pM Ha:pF≠pMHa:pF≠pM

None of these
Answer #1

2.

Given that,
sample one, x1 =12, n1 =40, p1= x1/n1=0.3
sample two, x2 =22, n2 =78, p2= x2/n2=0.282
null, Ho: p1 = p2
alternate, H1: p1 > p2
level of significance, α = 0.1
from standard normal table,right tailed z α/2 =1.282
since our test is right-tailed
reject Ho, if zo > 1.282
we use test statistic (z) = (p1-p2)/√(p^q^(1/n1+1/n2))
zo =(0.3-0.282)/sqrt((0.288*0.712(1/40+1/78))
zo =0.204
| zo | =0.204
critical value
the value of |z α| at los 0.1% is 1.282
we got |zo| =0.204 & | z α | =1.282
make decision
hence value of |zo | < | z α | and here we do not reject Ho
p-value: right tail - Ha : ( p > 0.2038 ) = 0.41926
hence value of p0.1 < 0.41926,here we do not reject Ho
ANSWERS
---------------
null, Ho: p1 = p2
alternate, H1: p1 > p2
test statistic: 0.204
critical value: 1.282
decision: do not reject Ho
p-value: 0.41926
we do not have enough evidence to support the claim that young men more likely to experience ADD/ADHD symptoms than young women